d2x/dt2=x-2y, d2y/dt2=4x+5y find general solution using matrix method.


1
Expert's answer
2022-02-08T12:22:00-0500

"z''=\begin{pmatrix}
 1 & -2 \\
 4 & 5
\end{pmatrix}z"


"\begin{vmatrix}
 1-\lambda & -2 \\
 4 & 5-\lambda
\end{vmatrix}=0"


"(1-\lambda)(5-\lambda)+8=0"

"\lambda^2+6\lambda -13=0"

"\lambda=\frac{-6\pm \sqrt{36+52}}{2}=-3\pm \sqrt{22}"


for "\lambda_1=-3-\sqrt{22}" :

"(4+\sqrt{22})x-2y=0"

"4x+(8+\sqrt{22})y=0"

"x\sqrt{22}-(10+\sqrt{22})y=0"

eigenvector:

"\begin{pmatrix}
 1 \\
 1+10/\sqrt{22}
\end{pmatrix}"


for "\lambda_2=-3+\sqrt{22}" :

"(4-\sqrt{22})x-2y=0"

"4x+(8-\sqrt{22})y=0"

"-x\sqrt{22}-(10-\sqrt{22})y=0"

eigenvector:

"\begin{pmatrix}
 1 \\
 1-10/\sqrt{22}
\end{pmatrix}"


"z'(t)=c_1e^{-(3+\sqrt{22})t}\begin{pmatrix}
 1 \\
 1+10/\sqrt{22}
\end{pmatrix}+c_2e^{-(3-\sqrt{22})t}\begin{pmatrix}
 1 \\
 1-10/\sqrt{22}
\end{pmatrix}"


"z(t)=C_1e^{-(3+\sqrt{22})t}\begin{pmatrix}
 1 \\
 1+10/\sqrt{22}
\end{pmatrix}+C_2e^{-(3-\sqrt{22})t}\begin{pmatrix}
 1 \\
 1-10/\sqrt{22}
\end{pmatrix}"